# TwinBrainVLA[Uncaptioned image]: Unleashing the Potential of Generalist VLMs for Embodied Tasks via Asymmetric Mixture-of-Transformers

### v1

@@ 4.2 Flow-Matching Action Expert
To enable high-precision, continuous control, we move beyond the discrete tokenization paradigm (Kim et al., 2024) and adopt the state-of-the-art generative policy architecture utilized in recent foundation models like Isaac-GR00T (NVIDIA et al., 2025) and $\pi_{0}$ (Intelligence et al., 2025). Specifically, our Action Expert employs the Diffusion Transformer (DiT) architecture (Peebles and Xie, 2023) and is trained via flow matching.
Architecture and Conditioning. Following the design in Isaac-GR00T, the Action Expert operates as a conditional decoder that denoises action trajectories. It takes a noisy action sequence $\mathbf{a}_{t}$ and the current timestep $t$ as input. While the policy architecture aligns with established baselines, a key distinction in our framework lies in the source of the condition. We inject the spatially-rich representations from our trainable Right Brain, $\mathbf{H}_{R}$, into the DiT via cross-attention layers. This ensures that the proven control capabilities of the DiT are guided explicitly by the embodied perception features extracted by our asymmetric backbone.
Flow-Matching Formulation. We utilize the flow-matching objective to train the policy, as demonstrated effectively in prior work (NVIDIA et al., 2025). The process models a conditional probability path flowing from a standard Gaussian prior $\mathbf{a}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to the ground-truth action distribution $\mathbf{a}_{1}$. The vector field regresson loss is defined as:
@@
where $v_{\psi}$ is the DiT network, $t\sim\mathcal{U}[0,1]$, and the target vector field is the straight line $\mathbf{a}_{1}-\mathbf{a}_{0}$. During inference, actions are synthesized by solving the Ordinary Differential Equation (ODE) using an Euler solver, translating the Right Brain’s understanding into smooth robotic movements.
@@ 4.3 Training Strategy
Optimization Objective. Consistent with standard VLA fine-tuning paradigms (NVIDIA et al., 2025), we train TwinBrainVLA using exclusively the robotic action objective. We do not incorporate auxiliary Next-Token Prediction (NTP) losses or mix in general vision-language datasets during this stage. The training objective is to minimize the Flow-Matching loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{FM}}(\theta_{R},\psi,\phi;\mathcal{D}_{\text{robot}}),$ (9)
@@
Asymmetric Update Rule. We enforce a strict parameter update policy. Let $\theta_{L}$ be the parameters of the Left Brain. During backpropagation, we set $\nabla\theta_{L}=0$. Gradients originating from the Action Expert flow backwards through the DiT ($\psi$), permeate the Right Brain ($\theta_{R}$) via the conditioning signals, and update the State Encoder ($\phi$). Crucially, at the AsyMoT fusion layer, the gradient flow is explicitly blocked (stop-gradient) at the Key-Value interface from the Left Brain. This ensures that the Left Brain acts as a stable semantic anchor, providing consistent features $K_{L},V_{L}$ without its weights being perturbed by the high-variance gradients typical of robotic control tasks.
@@ 5 Experiment
5.1 Experiments on SimplerEnv
Implementation and Training Setup. To demonstrate the scalability and effectiveness of our architecture, we instantiate TwinBrainVLA with two state-of-the-art VLM backbones: Qwen2.5-VL-3B-Instruct and Qwen3-VL-4B-Instruct. Consistent with our asymmetric design, both the frozen Left Brain and the trainable Right Brain are initialized from these pre-trained checkpoints to ensure aligned feature spaces. For the training data, we utilize two large-scale subsets from the Open X-Embodiment (OXE) dataset: the Bridge-V2 and the Fractal dataset. We are actively working on scaling our training to the complete OXE dataset, and future experimental results will be reported in subsequent updates. Comprehensive implementation details are provided in Appendix A.
Evaluation. The benchmark consists of four manipulation tasks: "put spoon on towel", "put carrot on plate", "stack green block on yellow block", "put eggplant in the yellow basket". For each task, we evaluate our VLA policy using the evaluation script provided by the SimplerEnv repository (Li et al., 2024c). To mitigate randomness, we run 480 independent trials and report the mean performance (Avg@480).
Baselines. We benchmark TwinBrainVLA against a comprehensive suite of state-of-the-art VLA policies, including RT-1-X (O’Neill et al., 2024), Octo (Team et al., 2024), OpenVLA (Kim et al., 2024), RoboVLM (Li et al., 2024b), TraceVLA (Zheng et al., 2025c), SpatialVLA (Qu et al., 2025b), CogACT (Li et al., 2024a), VideoVLA (Shen et al., 2025), $\pi_{0}$ (Black et al., 2024), and $\pi_{0.5}$ (Intelligence et al., 2025). All baseline results are obtained from their original papers or official reproductions to ensure a fair and reliable comparison.
@@ 5.1 Experiments on SimplerEnv
Table 1: Results of evaluating the VLA models with the WidowX robot in the SimplerEnv simulation environment. We highlight the best results in bold and the second-best results with underline.
@@
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1 Octo-Base (Team et al., 2024) 15.8 12.5 0.0 41.7 17.5 Octo-Small (Team et al., 2024) 41.7 8.2 0.0 56.7 26.7 OpenVLA (Kim et al., 2024) 4.2 0.0 0.0 12.5 4.2 OpenVLA-OFT (Kim et al., 2025) 12.5 4.2 4.2 72.5 23.4 RoboVLM (Li et al., 2024b) 50.0 37.5 0.0 83.3 42.7 TraceVLA (Zheng et al., 2025c) 12.5 16.6 16.6 65.0 27.7 SpatialVLA (Qu et al., 2025b) 20.8 20.8 25.0 70.8 34.4 CogACT (Li et al., 2024a) 71.7 50.8 15.0 67.5 51.3 VideoVLA (Shen et al., 2025) 75.0 20.8 45.8 70.8 53.1 $\pi_{0}$ (Black et al., 2024) 29.1 0.0 16.6 62.5 27.1 $\pi_{0.5}$ (Intelligence et al., 2025) 49.3 64.7 44.7 69.7 57.1 QwenGR00T + Qwen3-VL-3B (starVLA, 2025) 87.5 50.0 29.2 54.2 55.2 QwenGR00T + Qwen3-VL-8B (starVLA, 2025) 70.8 38.9 22.2 65.3 49.3 QwenGR00T + Qwen2.5-VL-7B (starVLA, 2025) 59.2 30.8 3.3 44.2 34.4 Isaac-GR00T-N1.6-Bridge (Team et al., 2025b) 64.5 65.5 5.5 93.0 57.1 TwinBrainVLA + Qwen2.5-VL-3B-Instruct 83.3 41.7 31.5 77.1 58.4 TwinBrainVLA + Qwen3-VL-4B-Instruct 73.0 58.3 33.3 83.3 62.0
@@
Results. The quantitative results on the SimplerEnv benchmark are presented in Table 1. Notably, despite not undergoing large-scale pre-training for robotic action prediction, TwinBrainVLA achieves state-of-the-art performance among all listed methods. Our framework demonstrates strong generalizability across different VLM families, attaining competitive success rates of 58.4% with Qwen2.5-VL-3B-Instruct and 62.0% with Qwen3-VL-4B-Instruct. The latter configuration surpasses the strongest baseline, Isaac-GR00T-N1.6 (57.1%), by a notable margin of +4.9%, validating the effectiveness of our asymmetric dual-brain architecture in bridging high-level semantic understanding and low-level robotic control.
@@ 5.2 Experiments on RoboCasa
Evaluation. We evaluate our policy on the RoboCasa GR1 Tabletop Benchmark, which comprises a diverse suite of 24 tabletop manipulation tasks. These tasks involve complex interactions with articulated objects and varying geometries, such as PnPBottleToCabinetClose (placing a bottle into a cabinet and closing the door), PnPCanToDrawerClose, and various microwave/toaster interaction scenarios. To ensure statistical reliability while maintaining consistency with our previous experiments, we report the success rate averaged over multiple independent trials (Avg@50) for each of the 24 tasks.

### v2

@@ 4.2 Flow-Matching Action Expert
Following the paradigm of recent mainstream VLAs (NVIDIA et al., 2025; Intelligence et al., 2025), we employ a generative action expert based on flow matching to enable high-precision continuous control. This module is implemented as a Diffusion Transformer (DiT) (Peebles and Xie, 2023), which operates as a conditional decoder. Specifically, it takes the representations $\mathbf{H}_{R}$ from the Right Brain as the condition to synthesize continuous action trajectories from noise. The training objective is defined by the standard vector field regression loss:
@@
where $v_{\psi}$ denotes the DiT network, $\mathbf{a}_{0}$ is the Gaussian noise, and $\mathbf{a}_{1}$ represents the ground-truth action.
@@ 4.3 Training Strategy
Optimization Objective. Consistent with standard VLA fine-tuning paradigms (NVIDIA et al., 2025), we train TwinBrainVLA using exclusively the robotic action objective. The training objective is to minimize the Flow-Matching loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{action}}(\theta_{R},\psi,\phi;\mathcal{D}_{\text{robot}}),$ (8)
@@ 5 Experiment
5.1 Experiments on SimplerEnv (OOD)
@@
Table 1: Results of evaluating the VLA models with the WidowX robot in the SimplerEnv simulation environment. We highlight the best results in bold and the second-best results with underline. Here, Vanilla VLA refers to a VLA constructed with a single VLM backbone. Its only difference from TwinBrainVLA is the removal of the frozen Left Brain.
@@
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1 Octo-Base (Team et al., 2024) 15.8 12.5 0.0 41.7 17.5 Octo-Small (Team et al., 2024) 41.7 8.2 0.0 56.7 26.7 OpenVLA (Kim et al., 2024) 4.2 0.0 0.0 12.5 4.2 RoboVLM (Li et al., 2024b) 50.0 37.5 0.0 83.3 42.7 TraceVLA (Zheng et al., 2025c) 12.5 16.6 16.6 65.0 27.7 SpatialVLA (Qu et al., 2025) 20.8 20.8 25.0 70.8 34.4 ThinkAct (Huang et al., 2025) 58.3 37.5 8.7 70.8 CogACT (Li et al., 2024a) 71.7 50.8 15.0 67.5 51.3 VideoVLA (Shen et al., 2025) 75.0 20.8 45.8 70.8 53.1 $\pi_{0}$ (Black et al., 2024) 29.1 0.0 16.6 62.5 27.1 $\pi_{0.5}$ (Intelligence et al., 2025) 49.3 64.7 44.7 69.7 57.1 Isaac-GR00T-N1.6-Bridge (Team et al., 2025b) 64.5 65.5 5.5 93.0 57.1 Vanilla VLA + Qwen2.5-VL-3B-Instruct 74.9 44.7 9.3 81.2 52.5 Vanilla VLA + Qwen2.5-VL-7B-Instruct 68.7 35.4 16.7 62.5 44.8 Vanilla VLA + Qwen3-VL-4B-Instruct 87.5 50.0 29.2 54.2 55.2 Vanilla VLA + Qwen3-VL-8B-Instruct 68.7 38.5 30.2 87.9 56.3 ours TwinBrainVLA + Qwen2.5-VL-3B-Instruct 83.3 41.7 31.5 77.1 58.4 TwinBrainVLA + Qwen3-VL-4B-Instruct 87.5 58.3 33.3 79.1 64.5
Implementation and Training Setup. To demonstrate the scalability and effectiveness of our architecture, we instantiate TwinBrainVLA with two state-of-the-art VLM backbones: Qwen2.5-VL-3B-Instruct and Qwen3-VL-4B-Instruct. Consistent with our asymmetric design, both the frozen Left Brain and the trainable Right Brain are initialized from these pre-trained checkpoints to ensure aligned feature spaces. For the training data, we utilize two large-scale subsets from the Open X-Embodiment (OXE) dataset: the Bridge-V2 and the Fractal dataset. Comprehensive implementation details are provided in Appendix B.
Evaluation. The benchmark consists of four manipulation tasks: "put spoon on towel", "put carrot on plate", "stack green block on yellow block", "put eggplant in the yellow basket". For each task, we evaluate our VLA policy using the evaluation script provided by the SimplerEnv repository (Li et al., 2024c). To mitigate randomness, we run 480 independent trials and report the mean performance (Avg@480). Since the evaluation environments are completely absent from the training set, this constitutes an Out-of-Domain (OOD) evaluation.
@@
Results. The quantitative results on the SimplerEnv benchmark are presented in Table 1. Notably, despite not undergoing large-scale pre-training for robotic action prediction, TwinBrainVLA achieves state-of-the-art performance among all listed methods. Our framework demonstrates strong generalizability across different VLM families, attaining competitive success rates of 58.4% with Qwen2.5-VL-3B-Instruct and 64.5% with Qwen3-VL-4B-Instruct. The latter configuration surpasses the strongest baseline, Isaac-GR00T-N1.6 (57.1%), by a notable margin of +7.4%, validating the effectiveness of our asymmetric dual-brain architecture in bridging high-level semantic understanding and low-level robotic control.
@@ 5.2 Experiments on RoboCasa
Table 2: Results of evaluating the VLA models with the GR1 robot in the RoboCasa Tabletop simulation environment. The results for Isaac-GR00T N1.5 and Isaac-GR00T N1.6 are sourced from the official Isaac-GR00T github repository (Team et al., 2025b). The results for the first four baseline methods are sourced from the official starVLA experiments (starVLA, 2025). Performance on all 24 tasks can be found in Table 8 of the Appendix.
Model QwenFAST +Qwen3VL QwenGR00T +Qwen3VL QwenPI +Qwen3VL QwenOFT +Qwen3VL Average 39.0 47.8 43.9 48.8 Model Isaac-GR00T N1.5 Isaac-GR00T N1.6 TwinBrainVLA + Qwen2.5VL TwinBrainVLA + Qwen3VL Average 48.2 47.6 53.5 54.6
@@
Evaluation. We evaluate our policy on the RoboCasa GR1 Tabletop Benchmark, which comprises a diverse suite of 24 tabletop manipulation tasks. These tasks involve complex interactions with articulated objects and varying geometries, such as PnPBottleToCabinetClose (placing a bottle into a cabinet and closing the door) and PnPCanToDrawerClose. To ensure statistical reliability while maintaining consistency with our previous experiments, we report the success rate averaged over 50 independent trials for each of the 24 tasks.
5.3 Experiments on LIBERO Benchmark
We also evaluate TwinBrainVLA on the LIBERO benchmark (Liu et al., 2023a). However, given that current VLA research has saturated this benchmark (over 95%) and its tasks are strictly in-domain, it offers limited insight into the benefits of general semantic preservation. To better evaluate the model’s generalization capabilities, we combined the training data from all four LIBERO task suites for VLA training and evaluated a single model across all four suites. In this setting, TwinBrainVLA attains a 97.6% average success rate without relying on massive pre-training. See Appendix F for implementation and evaluation details.
Table 3: Comparison on the LIBERO benchmark.
Method Spatial Object Goal Long Avg
OpenVLA 87.4 88.4 79.2 53.7 76.5
$\pi_{0}$ 96.8 98.8 95.8 85.2 94.1
$\pi_{0.5}$ 98.8 98.2 98.0 92.4 96.9
TwinBrainVLA 99.2 99.0 96.8 95.4 97.6
5.4 Real-robot Experiments
To demonstrate the practical applicability of TwinBrainVLA in real world, we conduct experiments using a Franka Research 3 (7-DOF).
Figure 4: Real-robot Experiment Environment.
Pre-training Setup. We curate a high-quality pre-training corpus by selecting 20 subsets from the Open X-Embodiment (OXE) dataset. Our selection criteria prioritize high video resolution, diverse task instructions, and unified end-effector position control formats. This dataset covers multiple robot embodiments, including WidowX, Google Robot, and Franka (see Appendix A for details). TwinBrainVLA is pre-trained on this datasets to acquire generalizable sensorimotor priors before specific task adaptation.
Fine-tuning and Evaluation. For downstream adaptation, we collect 300 teleoperated demonstration trajectories on the Franka robot for a pick-and-place task (instruction: “pick up the red block and put it into the gray box”). To rigorously assess robustness, we design three evaluation protocols: (1) In-Domain: The test environment and objects are identical to the training data; (2) Out-of-Domain (OOD): The target object’s attributes (e.g., block color) are altered to evaluate zero-shot visual generalization. (3) Pick-All: In this task, the robot is required to place all blocks into the box. We expect the model to complete this long-horizon task by leveraging the atomic actions learned from the training set. We report the average success rate over 30 independent trials for each setting.
Results. In real-robot experiments, despite utilizing significantly less pre-training data than $\pi_{0.5}$, our method achieves comparable performance after fine-tuning on teleoperation data. Furthermore, it demonstrates robust results in both OOD and Pick-All evaluations. This validates the effectiveness of our model architecture and underscores the importance of retaining generalizability for task generalization.
Table 4: Real-Robot Experiment Results. Success rates (%) of the pick-and-place task on the Franka Research 3 robot.
Method Success Rate
In-Domain Out-of-Domain Pick-All
OpenVLA (Kim et al., 2024) 7/30 0/30 0/30
$\pi_{0}$ (Black et al., 2024) 23/30 9/30 0/30
$\pi_{0.5}$ (Intelligence et al., 2025) 28/30 13/30 2/30
Vanilla VLA 22/30 9/30 0/30
TwinBrainVLA (Ours) 28/30 15/30 3/30
5.5 Ablation Experiments
Impact of asymmetric freezing strategy. To validate our asymmetric design, we ablate the freezing strategy by making the Left Brain trainable. Results show that this configuration hurts performance, causing a 7% drop in success rate for the Qwen3-VL-4B backbone on SimplerEnv. This result highlights that explicitly retaining general capabilities as an accessible resource is paramount for achieving high performance in VLA tasks.
Table 5: Ablation on Freezing Strategy. Quantitative evaluation of success rates (%) on SimplerEnv tasks. We benchmark TwinBrainVLA against the Single-Stream VLA (QwenGR00T) under identical training settings using Qwen2.5-VL-3B and Qwen3-VL-4B backbones.
Method Put Spoon on Towel Put Carrot on Plate Stack Block Eggplant in Basket Avg Qwen2.5-VL-3B-Instruct No Freezing 64.6 45.8 20.8 64.6 49.0 TwinBrainVLA 83.3 41.7 31.5 77.1 58.4 Qwen3-VL-4B-Instruct No Freezing 85.4 50.0 29.1 70.8 58.8 TwinBrainVLA 87.5 58.3 33.3 79.1 64.5
Impact of interaction density in AsyMoT. In TwinBrainVLA, joint attention is performed layer-wise by default. We further investigated the impact of interaction density by sparsifying the connections. Specifically, we experimented with interaction intervals of $k=\{0,1,2,4,8\}$ layers. When $k=1$, the asymmetric joint attention and naive self-attention are performed alternately. The results in Table 6 demonstrate that effective information fusion yields performance gains, whereas performance degrades when the connections become overly sparse.
Impact of dual-VLM. If the left brain and the attention connections between the two VLMs are removed, the network architecture degenerates into a Vanilla VLA. As shown in Table 1, this degradation leads to a performance drop of nearly 7 percentage points.
Table 6: Ablation on AsyMoT Interaction Frequency. We investigate the impact of sparsifying the joint attention connections by varying the interaction interval $k$ on SimplerEnv benchmark.
Interval ($k$) 0 (ours) 1 2 4 8 Qwen2.5-VL-3B-Instruct based 58.4 59.4 57.4 56.5 55.9 Qwen3-VL-4B-Instruct based 64.5 64.1 66.7 62.5 61.9
6 Discussion
6.1 From Forgetting Mitigation to Capability Utilization
TwinBrainVLA effectively prevents catastrophic forgetting by maintaining a strictly frozen "Left Brain" throughout the training process. This strategy explicitly preserves the general semantic understanding capabilities acquired during VLM pre-training, ensuring that the model retains its robust open-world knowledge while the "Right Brain" specializes in embodied control.
Figure 5: TwinBrainVLA Attention Map Visualization at Runtime. It can be seen that the right brain attends to the semantic information of the left brain via the AsyMoT mechanism.
6.2 The Necessity of General Semantics for Embodied Intelligence
The fundamental premise of VLA models is to harness the extensive general capabilities acquired during the pre-training of VLMs to achieve generalized embodied intelligence. However, current VLA training paradigms often suffer from severe overfitting to specific robotic datasets (Wanna et al., 2026), leading to "catastrophic forgetting" of the backbone’s original capabilities. Addressing this paradox is the primary motivation behind TwinBrainVLA. TwinBrainVLA addresses this by explicitly preserving the backbone’s semantic integrity. Our results confirm that retaining general capabilities directly boosts performance: our approach matches commercial models (e.g., $\pi_{0.5}$) without massive pre-training. Furthermore, this preserved knowledge enables robust instruction following and semantic generalization in real-world scenarios, handling complex tasks that traditional fine-tuning struggles to address."
6.3 "Twin-to-One" via Knowledge Distillation
While TwinBrainVLA successfully achieves the utilization of general capabilities by incorporating a frozen VLM as a semantic anchor, we acknowledge that this dual-stream architecture inevitably introduces additional computational overhead during deployment. To bridge the gap between high-level capability utilization and low-latency deployment, we conducted a preliminary exploration into a "Twin-to-One" knowledge distillation paradigm. In this experiment, we employed the feature representations from the Right Brain as a source of supervisory signals to re-train a Vanilla VLA architecture. The results revealed a tangible improvement in performance, which substantiates that TwinBrainVLA achieves superior training outcomes through the effective fusion of the Left Brain. This finding also highlights the potential for deeper exploration into internalizing these fused capabilities into simpler architectures. The details of these experiments are provided in Appendix G.
7 Conclusion
In this work, we identified a critical paradox in current VLA models: the fine-tuning process required for robotic control inevitably leads to the "catastrophic forgetting" of the VLM backbone’s general capabilities. In order to harness the pre-trained general capabilities of VLMs to achieve generalization in embodied AI, we introduce TwinBrainVLA[Uncaptioned image] which is a framework that structurally decouples multimodal understanding from sensorimotor control. By orchestrating a frozen "Left Brain" for open-world understanding and a trainable "Right Brain" for embodied action via Asymmetric Mixture-of-Transformers (AsyMoT) mechanism, our architecture allows the policy to query high-level semantics without corrupting them. Extensive empirical evaluations on SimplerEnv, RoboCasa benchmarks and real-robot experiments demonstrate that TwinBrainVLA significantly outperforms mainstream VLAs.
While TwinBrainVLA marks a meaningful step towards general embodied AI, the challenges of achieving higher efficiency and deeper semantic fusion remain. We provide a detailed discussion on these limitations and future directions in Appendix H.
Appendix A Real Robot Experiments Detail
This section describes the training implementation and evaluation setup for our real-robot experiments.
A.1 Real Robot Pre-training
Table 7 presents the OXE subsets used for our pre-training.
Table 7: Selected OXE Subsets for Pre-training. We curate 20 subsets from the Open X-Embodiment dataset that utilize end-effector position control. The table details the dataset source, the robot platform, and the number of episodes utilized.
Dataset Name Robot Episodes
Berkeley Bridge WidowX 25,460
RT-1 Robot Action Google Robot 79,499
DROID Franka 92,233
BC-Z Google Robot 39,350
Stanford HYDRA Franka 550
Austin VIOLA Franka 135
Berkeley Autolab UR5 UR5 896
NYU Franka Play Franka 456
Stanford Kuka Multimodal Kuka iiwa 3,000
Freiburg Franka Play Franka 3,242
USC Jaco Play Jaco 2 976
NYU VINN Hello Stretch 435
Austin BUDS Franka 50
UCSD Kitchen xArm 150
CMU Franka Pick-Insert Data Franka 520
Austin Mutex Franka 1,500
Berkeley Fanuc Manipulation Fanuc Mate 415
CMU Play Fusion Franka 576
CMU Stretch Hello Stretch 135
DobbE Hello Stretch 5,208
Total 254,786
A.2 Data Collection via Teleoperation
We use Polymetis (Lin et al., 2021) for teleoperation and collect human expert data to support imitation learning for VLA models. For real-world experiments, we utilize a Franka Research 3 robot arm equipped with two Intel RealSense D435 cameras. We collected 300 expert human demonstrations, following the instruction format: “pick up the {color} block and put it into the gray box”. The dataset is balanced, containing 100 trajectories for each of the three colors: red, green, and blue. A schematic illustration of the demonstration samples is presented in Figure 6.
Figure 6: Diagram of human expert data acquired through teleoperation
A.3 Fine Tuning on Human Expert Demonstration Samples
We perform fine-tuning on the checkpoints pretrained on the OXE subset, using the human expert demonstration data we collected. The training is conducted on 8 $\times$ H100 GPUs with a per-device batch size of 16 for 80K steps.
A.4 Real Robot Evaluation
Our evaluation is divided into three settings: in-domain, out-of-domain and pick-all.
  • In-Domain. In this setting, the evaluation tasks use exactly the same objects and task instructions as those in the training set, with the positions of the blocks altered randomly.
  • Out-of-Domain. In this setting, the instructions for the evaluation tasks differ from those in the training set, primarily in terms of the color of the blocks to be grasped being altered. Specifically, the OOD evaluation task is defined by the instruction: “pick up the yellow block and put it into the gray box”.
  • Pick-All. In this setting, the task instruction is distinct from those in the training set, requiring the robotic arm to place all blocks into the box. Specifically, the task is defined by the instruction: “pick up all blocks and put them into the gray box“, with four blocks (red, blue, green, and yellow) present in the workspace. The model is required to comprehend the task semantics and achieve long-horizon generalization by composing atomic actions learned from the training dataset.
Appendix B Implementation Details
For fine-tuning, we set the default training duration to 40K steps. This choice aligns with standard practices in related work and ensures that the training can be completed within one day on NVIDIA H100 GPUs.
Our model architecture incorporates a state encoder designed to process proprioceptive information, such as joint angles and gripper states. However, since mainstream simulation benchmarks (e.g., SimplerEnv) primarily evaluate vision-only policies and do not typically involve state observations in their standard protocols, we omitted the state encoder during all simulation experiments to ensure fair comparison. Conversely, in real-robot settings where precise proprioception is crucial for control, the state encoder was active and utilized throughout both the pre-training and fine-tuning phases.
Figure 7: Examples from the Bridge V2 dataset in the training set.
Appendix C QwenGR00T, QwenPI and QwenOFT
For certain simulation experiments, we adopted the built-in baselines provided by StarVLA, re-implementing, training, and evaluating them under our uniform experimental setup. The architectures of these baselines are detailed in this section. As illustrated in Figure 8, these models differ primarily in their action modeling and VLM-policy interaction mechanisms:
  • QwenGR00T (Figure 8a): Incorporating a Flow Matching-based Action Expert, this model follows a late-fusion strategy (NVIDIA et al., 2025). The VLM functions as a high-level semantic encoder, passing only its last hidden states to the Action Expert. This design treats the VLM representation as a fixed holistic condition for the flow generation process.
  • QwenPI (Figure 8b): Also leveraging Flow Matching for continuous control, QwenPI adopts a deep-fusion architecture (Black et al., 2024). Unlike GR00T, it establishes dense connections between the VLM and the Action Expert, allowing for layer-wise attention interactions. This enables the policy to access and leverage multi-scale semantic features from the VLM’s intermediate layers during action generation.
  • QwenOFT (Figure 8c): Representing the discrete control paradigm (Kim et al., 2024), this model discretizes the action space into token bins and predicts actions autoregressively, treating manipulation as a standard next-token prediction task.
The implementations of the aforementioned architectures are all adopted from the built-in modules of the starVLA (starVLA, 2025) framework and are not claimed as our original contributions.
Figure 8: Architectural overview of VLA baselines with distinct fusion strategies. (a) QwenGR00T adopts a late-fusion design where a flow-matching Action Expert is conditioned solely on the last hidden states of the VLM. (b) QwenPI employs a deep-fusion strategy, enabling the flow-matching policy to interact with the VLM via layer-wise attention mechanisms. (c) QwenOFT follows the discrete token paradigm, predicting action bins autoregressively.
Appendix D RoboCasa Tabletop Benchmark
The RoboCasa Tabletop benchmark comprises 24 distinct sub-tasks. We present the detailed success rates for all these tasks in Table 8.
Figure 9: Schematic of RoboCasa Tabletop Evaluation.